If 10 cubic feet of water weigh approximatley 625 pounds, what is the weight of the water in a box shaped container, measuring 3 yards by 5 yards by 2 feet, that has been filled to capacity with water? Thanks
Answer by checkley75(3666)

9*15*2=270 CUBIC FEET OF WATER.
270*625/10=270*62.5=16,875 POUNDS.